# Unsteady aerodynamic theory and experiments of hovering membrane wings

### v1

@@ 1 Introduction
In a parallel theoretical effort to understand the unsteady aerodynamic performance of compliant membrane aerofoils, Tiomkin & Jaworski (2022) extended the classical unsteady thin aerofoil theory of von Kármán & Sears (1938) and Sears (1940), originally derived for a flat plate immersed in inviscid incompressible flow, to include the aeroelastic response of the membrane aerofoil to unsteady flow conditions due to a prescribed flapping motion or a transverse gust encounter. The extended theory provides closed-form expressions for the unsteady lift coefficient due to membrane deformation in inviscid incompressible unsteady flow assuming low angles of attack, small membrane camber, and a constant freestream velocity. The analytical expressions yield the unsteady lift coefficient for any prescribed membrane deformation in time, where the deformation can be obtained either by solving the membrane dynamic equation, as demonstrated by Tiomkin & Jaworski (2022), where simplifying assumptions such as constant tension along the membrane are generally required, or directly from experimental measurements.
@@ 2.1 Theoretical model
We consider a membrane aerofoil of thickness $h$ that is held by supports at a distance $c=2b$ from the leading to the trailing edge. The membrane is immersed in a uniform inviscid incompressible flow of fluid density $\rho$ and variable speed $U(t)$ at an angle of attack $\alpha$, as encountered by an aerofoil section along the span of a hovering membrane wing (see figure 1b). The unsteady lift acting on the membrane aerofoil is derived here assuming small membrane deformation $(|y_{x}|\ll 1)$ and attached flow at high angles of attack. These assumptions are justified by the experimental results of Gehrke & Mulleners (2025), in which unsteady membrane deformation was shown to suppress the formation of leading edge vortices and flow separation, even at angles of attack as high as $55^{\circ}$ for large parts of the flapping cycle (see figures 1d and 1f).
@@
in which $b,b/\overline{U}$ and $U(t)$ are used as the units of length, time, and circulation per unit length, respectively. Here $\gamma$ is the normalised bound vorticity per unit length along the profile, the dashed integral denotes the Cauchy principal value, and $\gamma_{w}(\eta,\xi)$ describes the normalised vorticity per unit length at location $\xi$ along the wake, $\xi\in\left[1,1+\eta\right]$, at time $t$ when the aerofoil has travelled a total distance of $\eta(t)$.
@@
By following the standard application of Söhngen’s inversion formula to (2) and enforcing Kelvin’s theorem (cf., Söhngen, 1939; Bisplinghoff et al., 1996, p. 289) we obtain
@@
where we apply the standard coordinate transformation, $x=-\cos\theta$, which places the aerofoil leading edge at x=1x=-1 ($\theta=0$) and the trailing edge at $x=1$ ($\upi\theta=\upi$). By substituting (6) into the normal velocity expression (1) and applying the Laplace transform to (4) we obtain an expression for the wake vorticity distribution in the Laplace domain,
@@
where $f_{m}(\eta)$ is the $\eta$-domain representation of (8). Note that $g_{m}(\eta)$ represents the normalised non-circulatory lift coefficient, while $f_{m}(\eta)$ describes the circulatory term, as it is directly related to the wake vorticity through (7). For a constant freestream $(\eta\equiv t)$ at small angles of attack $(\cos{\alpha}\cong 1)$ the solution described above in (11)-(13) coincides with the unsteady lift coefficient given by Tiomkin & Jaworski (2022).
@@
$C_{n_{d}}(t)=2\pi V(t)^{2}\left\{C(k){f}_{m}\left(\eta\left(t\right)\right)+g_{m}\left(\eta\left(t\right)\right)\right\},$ (14)
@@
$C_{l_{d}}(t)=C_{n_{d}}\cos{\alpha}=2\pi\cos{\alpha}\,V(t)^{2}\left\{C(k){f}_{m}\left(\eta\left(t\right)\right)+g_{m}\left(\eta\left(t\right)\right)\right\},$ (15)
@@ 2.2 Experimental configuration
This configuration enables passive deformation of the thin aerofoil shape under aerodynamic loading, creating a natural synergy between structural flexibility and aerodynamic forces, characterised by the aeroelastic number, $\textit{Ae}={Eh}/({\frac{1}{2}\rho\overline{U}^{2}c})$, where $E$ is the Young modulus of the membrane. The model wing performs a sinusoidal stroke along the horizontal plane, and a trapezoidal angle of attack ($\alpha$) profile (see figures 2b), similar to the hovering kinematics of small bats and insects. The experimental study focused on capturing the membrane dynamic response and the unsteady aerodynamic forces produced by the flapping wing for a wide range of aeroelastic numbers, $\textit{Ae}=0.25-12$, and angle-of-attack amplitudes of $\hat{\alpha}=15^{\circ}-75^{\circ}$. Aerodynamic forces and moments were recorded at the wing root with a 6-axis force/torque transducer, while the wing’s deformation was captured throughout the flapping cycle with two machine vision cameras in stereo-configuration and reconstructed using photogrammetry (figure 1b). Three-dimensional deformation results showed no significant spanwise variation in membrane profile, allowing its representation by a single two-dimensional profile (Gehrke & Mulleners, 2025). Finally, the unsteady flow field around the membrane wing was recorded using planar particle image velocimetry (PIV) at a fixed span-wise position, located at the radius of the second moment of area ($r=0.55R$, where $R$ is the total wingspan). More information about the experimental setup and in-depth aerodynamic analysis can be found in Gehrke et al. (2022) and Gehrke & Mulleners (2025).
@@ 2.3 Connecting theory to measurements
Figure 2: (a) Sketch of the flexible membrane wing profile, (b) Angle of attack ($\alpha$) and local wing velocity ($U$) profiles as a function of normalised time $t/T$, (c) Membrane camberline, $y(x)$, obtained at different times ($t/T=0$ to $0.5$), (d) Membrane Fourier coefficients, $\mathcal{F}_{n}$, as a function of time $t/T$, (e) Measured lift coefficients, $C_{L}$, as a function of time $t/T$ for a membrane and a rigid wing, (f) Finite-wing lift coefficient due to membrane deformation, $C_{L_{d}}$, as a function of time $t/T$; comparing the experimental measurements with the theory predictions. Very good agreement is observed between the theoretical and measured lift coefficients, computed for $\hat{\alpha}=35^{\circ},\textit{Ae}=2.5$.
Seeking to validate the unsteady aerodynamic theory derived in §2.1, membrane deformation measurements obtained by Gehrke et al. (2022) along the mid-wingspan section are converted first to body-fixed coordinates, normalised by the instantaneous semi-chord length, and decomposed into a Fourier series expansion (6) (see figures 2c and 2d, respectively). The resultant Fourier coefficients are then used to compute the theoretical sectional lift coefficient, $C_{l_{d}}(t)$, via (15). This result yields the finite-wing lift contribution from membrane deformation,
$C_{L_{d}}(t)=\frac{C_{L_{\alpha}}}{2\pi}\frac{c(t)}{\overline{c}}C_{l_{d}}(t),$ (17)
@@
The corresponding experimental value for the lift coefficient due to membrane deformation is obtained by calculating the difference between the lift coefficient measured for the membrane wing and the lift coefficient obtained on a rigid flat plate that undergoes an equivalent prescribed flapping motion (see figures 2e). The resultant experimental lift coefficient due to membrane deformation is then compared with the corresponding theoretical value (17) for every experimental scenario considered (e.g. figure 2f). Note that the gray areas in figures 2b and 2d-f mark the time intervals during which the angle of attack varies in time in the experiments, violating the theory assumptions. Therefore, the theoretical prediction, in which only effects of the instantaneous angle of attack are considered, is expected to agree with the experimental results for $0.08<t/T<0.42$, with a less favorable agreement in the gray-shaded areas.
@@ 3 Results and discussion
The unsteady lift coefficient of a finite membrane wing undergoing a prescribed flapping motion is studied for angle-of-attack amplitudes of $\hat{\alpha}=35^{\circ}$ and $55^{\circ}$ with various values of the aeroelastic number, as available from experimental data of Gehrke et al. (2022) (see figures 3i.a-c and 3ii.a-e, respectively). For each of the available scenarios, determined by $\hat{\alpha}$ and Ae, the experimental and theoretical lift coefficient results, and the maximum measured deformation are illustrated with blue-, red- and gray-shaded areas, respectively. The shaded regions show data from the forward and backward strokes, while the solid lines indicate their average. Since the wing flaps in a still-water tank, there is no imposed asymmetry between stroke directions. However, minor variations in membrane deformation arise between the two strokes due to experimental variability. Accordingly, we treat them as separate realizations of the same scenario (characterised by $\hat{\alpha}$ and Ae) when applying the aeroelastic model.
For an angle-of-attack amplitude of $\hat{\alpha}=35^{\circ}$ a very good agreement is observed between the theoretical prediction and the load-cell measurements obtained for moderate and high Ae values of $\textit{Ae}=2.5$ and $3.4$ (figures 3i.a and 3i.b, respectively). However, as the aeroelastic number is decreased to $1.7$, the maximum membrane camber increases to $26\%$ of the chord-length and the theory overestimates the lift coefficient amplitude compared to the measured value. This is typical of cases in which trailing-edge separation appears on highly cambered aerofoils, as is indeed seen in the PIV results of the low-Ae membrane (figure 3i.d). Notably, the general trend of the unsteady lift coefficient is captured well by the theory for all Ae values examined here for $\hat{\alpha}=35^{\circ}$, and is strongly correlated to the maximum membrane camber.
@@
Note that as we compare here results of the lift coefficient due to membrane deformation, the model assumptions need to hold with regards to the difference between the flowfield obtained for a rigid wing and the flowfield about a compliant membrane wing, where both wings perform the same kinematic motion. Therefore, cases in which the flowfield around the membrane wing is very close to the one obtained for the flexible wing (e.g. second row in 3i.d and 3ii.f depicting results for $\hat{\alpha}=35^{\circ},\textit{Ae}=5.2$ and $\hat{\alpha}=55^{\circ},\textit{Ae}=5$, respectively) produce very good agreement between the two methods.
For an angle-of-attack amplitude of $\hat{\alpha}=55^{\circ}$ we again see good agreement between the theoretical prediction and measurements of the unsteady lift coefficient due to membrane deformation obtained for aeroelastic numbers between 1.951.9-5, whereas values of $\textit{Ae}\leq 1.7$ present significant differences between theoretical and experimental results (figure 3ii.a-e). For the lowest value of $\textit{Ae}=0.95$ the theory overestimates the lift coefficient due to an excessive membrane camber of more than $30\%$, whereas for $\textit{Ae}=1.7$ a moderate maximum camber of $20\%$ is obtained and the predictive tool underestimates the lift coefficient. Flow field snapshots for $\textit{Ae}=0.85$ and $\textit{Ae}=1.9$ (figure 3ii.f, fourth and third lines, respectively) show that, within this range of the aeroelastic number, vortices are shed on the suction side near the trailing edge of the aerofoil. Accordingly, the theoretical model exhibits reduced accuracy at low aeroelastic numbers ($\textit{Ae}\leq 1.7$) due to the observed flow separation.

### v2

@@ 1 Introduction
In a parallel theoretical effort to understand the unsteady aerodynamic performance of compliant membrane aerofoils, Tiomkin & Jaworski (2022) extended the classical unsteady thin aerofoil theory of von Kármán & Sears (1938) and Sears (1940), originally derived for a flat plate immersed in inviscid incompressible flow, to include the aeroelastic response of the membrane aerofoil to unsteady flow conditions due to a prescribed flapping motion or a transverse gust encounter. The extended theory provides closed-form expressions for the unsteady lift coefficient due to membrane deformation in inviscid incompressible unsteady flow assuming low angles of attack, small membrane camber, and a constant freestream velocity. The analytical expressions yield the unsteady lift coefficient for any (small amplitude) membrane deformation in time, where the deformation can be obtained either by solving the membrane dynamic equation, as demonstrated by Tiomkin & Jaworski (2022), where simplifying assumptions such as constant tension along the membrane are generally required, or directly from experimental measurements.
@@ 2.1 Theoretical model
Figure 2: Representation of (a) the experimental membrane wing section and surrounding flowfield and (b) the theoretical model used to derive the unsteady lift coefficient acting on a flexible membrane aerofoil in variable freestream velocity. PIV results shown on the experimental model support the assumptions of attached flow and wake extension along the chord line, utilized in the theoretical model.
We consider a thin membrane aerofoil that is held by supports at a distance $c=2b$ from the leading to the trailing edge. The membrane is immersed in a uniform inviscid incompressible flow of fluid density $\rho$ and variable speed $U(t)$ at an angle of attack $\alpha$ (see figure 2b), as encountered by an aerofoil section along the span of a hovering membrane wing (figures 1b and 2a). The unsteady lift acting on the membrane aerofoil is derived here assuming small membrane deformation $(|y_{x}|\ll 1)$ and attached flow at high angles of attack. These assumptions are justified by the experimental results of Gehrke & Mulleners (2025), in which unsteady membrane deformation was shown to suppress the formation of leading edge vortices and flow separation, even at angles of attack as high as $55^{\circ}$ for large parts of the flapping cycle (see figures 1d and 1f).
@@
in which $b,b/\overline{U}$ and $U(t)$ are used as the units of length, time, and circulation per unit length, respectively. Here $\gamma$ is the normalised bound vorticity per unit length along the profile, the dashed integral denotes the Cauchy principal value, and $\gamma_{w}(\eta,\xi)$ describes the normalised vorticity per unit length at location $\xi$ along the wake, $\xi\in\left[1,1+\eta\right]$, at time $t$ when the aerofoil has travelled a total distance of $\eta(t)$ (see figure 2b).
@@
The above assumptions are supported by experimental observations of the wake (e.g. figure 2a), in which the wake extends along the aerofoil chord-line in the vicinity of the membrane trailing edge. By following the standard application of Söhngen’s inversion formula to (2) and enforcing Kelvin’s theorem (cf., Söhngen, 1939; Bisplinghoff et al., 1996, p. 289) we obtain
@@
where we apply the standard coordinate transformation, $x=-\cos\theta$, which places the aerofoil leading edge at x=1x=-1 ($\theta=0$) and the trailing edge at $x=1$ ($\upi\theta=\upi$). Note that the odd-index Fourier coefficients denote here the amplitude of the corresponding membrane mode-shape in the decomposition of the membrane deformation, while a combination of the even-index Fourier coefficients provides the even mode-shapes (see Tiomkin & Raveh, 2017). By substituting (6) into the normal velocity expression (1) and applying the Laplace transform to (4) we obtain an expression for the wake vorticity distribution in the Laplace domain,
@@
where $f_{m}(\eta)$ is the $\eta$-domain representation of (8). Note that $g_{m}(\eta)$ represents the normalised non-circulatory lift coefficient, while $f_{m}(\eta)$ describes the circulatory term, as it is directly related to the wake vorticity through (7). For a constant freestream at small angles of attack $(\cos{\alpha}\cong 1,\eta\equiv t)$ the solution described above in (11)-(13) coincides with the unsteady lift coefficient given by Tiomkin & Jaworski (2022).
@@
[…]
@@
[…]
@@
We further note that in cases where the first mode is significantly more dominant than all other modes $(\mathcal{F}_{1}\gg\mathcal{F}_{n}\;\forall\,n\neq 1)$, the membrane deformation can be represented by $y(\eta,\theta)=\frac{1}{2}\mathcal{F}_{1}(\eta)\sin^{2}{\theta}$, where $\frac{1}{2}\mathcal{F}_{1}(\eta)$ is the maximum camber of the aerofoil. Under these conditions a simplified expression is readily obtained for the lift coefficient that depends only on the wing kinematics and the first Fourier coefficient (or maximum membrane camber),
[…]
@@ 2.2 Experimental configuration
This configuration enables passive deformation of the thin aerofoil shape under aerodynamic loading, creating a natural synergy between structural flexibility and aerodynamic forces, characterised by the aeroelastic number, $\textit{Ae}={Eh}/({\frac{1}{2}\rho\overline{U}^{2}c})$, where $E$ is the Young modulus of the membrane. The model wing performs a sinusoidal stroke along the horizontal plane, and a trapezoidal angle of attack ($\alpha$) profile, similar to the hovering kinematics of small bats and insects (see figures 1b and 3b showing the experimental setup and wing kinematics). The experimental study focused on capturing the membrane dynamic response and the unsteady aerodynamic forces produced by the flapping wing for a wide range of aeroelastic numbers, $\textit{Ae}=0.25-12$, and angle-of-attack amplitudes of $\hat{\alpha}=15^{\circ}-75^{\circ}$. Aerodynamic forces and moments were recorded at the wing root with a 6-axis force/torque transducer, while the wing’s deformation was captured throughout the flapping cycle with two machine vision cameras in stereo-configuration and reconstructed using photogrammetry. Three-dimensional deformation results showed no significant spanwise variation in membrane profile, allowing its representation by a single two-dimensional profile (Gehrke & Mulleners, 2025). Finally, the unsteady flowfield around the membrane wing was recorded using planar particle image velocimetry (PIV) at a fixed span-wise position, located at the radius of the second moment of area ($r=0.55R$, where $R$ is the total wingspan). More information about the experimental setup, data acquisition and processing, and in-depth aerodynamic analysis can be found in Gehrke et al. (2022) and Gehrke & Mulleners (2025).
@@ 2.3 Connecting theory to measurements
Figure 3: (a) Sketch of the flexible membrane wing profile, (b) Angle of attack ($\alpha$) and local wing velocity ($U$) profiles as a function of normalised time $t/T$, (c) Membrane camberline, $y(x)$, obtained at different times ($t/T=0$ to $0.5$), (d) Membrane Fourier coefficients, $\mathcal{F}_{n}$, as a function of time $t/T$, (e) Measured lift coefficients, $C_{L}$, as a function of time $t/T$ for a membrane and a rigid wing, (f) Finite-wing lift coefficient due to membrane deformation, $C_{L_{d}}$, as a function of time $t/T$; comparing the experimental measurements with the theory prediction, along with simplified model predictions assuming single-mode oscillations ($\mathcal{F}_{1}$), constant freestream velocity ($\overline{U}$), and a quasi-steady aerodynamic model (QS). Very good agreement is observed between the theoretical and measured lift coefficients, computed for $\hat{\alpha}=35^{\circ},\textit{Ae}=2.5$, which is also captured by the single-mode approximation. However, the simplified constant $\overline{U}$ model and the quasi-steady model do not agree well with measurements.
Seeking to validate the unsteady aerodynamic theory derived in §2.1, membrane deformation measurements obtained by Gehrke et al. (2022) along the mid-wingspan section are converted first to body-fixed coordinates, normalised by the instantaneous semi-chord length, and decomposed into a Fourier series expansion (6) (see figures 3c and 3d, respectively). The resultant Fourier coefficients are then used to compute the theoretical sectional lift coefficient, $C_{l_{d}}(t)$, via (15). This result yields the finite-wing lift contribution from membrane deformation,
@@
where $c(t)$ and $\overline{c}$ are the instantaneous and reference chord-length, respectively. The finite-wing lift slope $C_{L_{\alpha}}$ is obtained using the two-station approximation (e.g. Anderson, 2024, pp. 451-454) for a rectangular wing of aspect ratio ÆR=2.7=2.7, matching the wing geometry used in the experiments of Gehrke et al. (2022).
The corresponding experimental value for the lift coefficient due to membrane deformation is obtained by calculating the difference between the lift coefficient measured for the membrane wing and the lift coefficient obtained on a rigid flat plate that undergoes an equivalent prescribed flapping motion (see figures 3e). The resultant experimental lift coefficient due to membrane deformation is then compared with the corresponding theoretical value (18) for every experimental scenario considered (e.g. figure 3f). Note that for cases of dominant first-mode oscillations the sectional lift coefficient in (18) can be obtained from the approximate expression (17). In the scenario presented in figure 3, the first Fourier coefficient is clearly the most dominant coefficient and therefore the lift coefficient predicted with the single mode approximation is practically indistinguishable from the full model prediction. However, for more general cases, in which higher mode oscillations are observable in the membrane dynamic response, the full model (15) is necessary to predict the sectional lift coefficient. Such higher-mode oscillations are expected, for example, in cases of higher reduced frequency, beyond the first natural frequency of the membrane. An alternative use of the standard unsteady aerodynamic model with constant freestream velocity significantly under-predicts the lift coefficient, while a quasi-steady model strongly over-predicts the response (figure 3f), highlighting the importance of including the effect of the variable freestream velocity and the shedding wake in the predictive model. Note that the gray areas in figures 3b and 3d-f mark the time intervals during which the angle of attack varies in time in the experiments, violating the theory assumptions. Therefore, the theoretical prediction, in which only effects of the instantaneous angle of attack are considered, is expected to agree with the experimental results for $0.08<t/T<0.42$, with a less favourable agreement in the gray-shaded areas.
@@ 3 Results and discussion
The unsteady lift coefficient of a finite membrane wing undergoing a prescribed flapping motion is studied for angle-of-attack amplitudes of $\hat{\alpha}=35^{\circ}$ and $55^{\circ}$ with various values of the aeroelastic number, as available from experimental data of Gehrke et al. (2022) (see figures 4i.a-c and 4ii.a-e, respectively). For each of the available scenarios, determined by $\hat{\alpha}$ and Ae, the experimental and theoretical lift coefficient results, and the maximum measured deformation are illustrated with blue-, red- and gray-shaded areas, respectively. The shaded regions show data from the forward and backward strokes, while the solid lines indicate their average. Since the wing flaps in a still-water tank, there is no imposed asymmetry between stroke directions. However, minor variations in membrane deformation arise between the two strokes due to experimental variability. Accordingly, we treat them as separate realizations of the same scenario (characterised by $\hat{\alpha}$ and Ae) when applying the aeroelastic model.
For an angle-of-attack amplitude of $\hat{\alpha}=35^{\circ}$ very good agreement is observed between the theoretical prediction and the load-cell measurements obtained for moderate and high Ae values of $\textit{Ae}=3.4$ and $2.5$ (figures 4i.a and 4i.b, respectively). However, as the aeroelastic number is decreased to $1.7$, the maximum membrane camber increases to $26\%$ of the chord-length at mid-stroke and the theory overestimates the lift coefficient amplitude compared to the measured value. This is typical of cases in which trailing-edge separation appears on highly cambered aerofoils, as is indeed seen in the PIV results of the low-Ae membrane (figure 4i.d). Notably, the general trend of the unsteady lift coefficient is captured well by the theory for all Ae values examined here for $\hat{\alpha}=35^{\circ}$, and is strongly correlated to the maximum membrane camber.
@@
Note that as we compare here results of the lift coefficient due to membrane deformation, the model assumptions need to hold with regards to the difference between the flowfield obtained for a rigid wing and the flowfield about a compliant membrane wing, where both wings perform the same kinematic motion. Therefore, cases in which the flowfield around the membrane wing is very close to the one obtained for the rigid wing (e.g. second row in 4i.d and 4ii.f depicting results for $\hat{\alpha}=35^{\circ},\textit{Ae}=5.2$ and $\hat{\alpha}=55^{\circ},\textit{Ae}=5$, respectively) produce very good agreement between the two methods.
For an angle-of-attack amplitude of $\hat{\alpha}=55^{\circ}$ we again see good agreement between the theoretical prediction and measurements of the unsteady lift coefficient due to membrane deformation obtained for aeroelastic numbers between 1.951.9-5, whereas values of $\textit{Ae}\leq 1.7$ present significant differences between theoretical and experimental results (figure 4ii.a-e). For the lowest value of $\textit{Ae}=0.95$ the theory overestimates the lift coefficient due to an excessive membrane camber of more than $30\%$, whereas for $\textit{Ae}=1.7$ a moderate maximum camber of $20\%$ is obtained and the predictive tool underestimates the lift coefficient. Flowfield snapshots for $\textit{Ae}=0.85$ and $\textit{Ae}=1.9$ (figure 4ii.f, fourth and third lines, respectively) show that, within this range of the aeroelastic number, vortices are shed on the suction side near the trailing edge of the aerofoil, remaining attached for $\textit{Ae}=1.9$} and detaching for lower Ae. Accordingly, the theoretical model exhibits reduced accuracy at low aeroelastic numbers ($\textit{Ae}\leq 1.7$) due to the observed flow separation.
The time-averaged error between the predicted and measured lift coefficient due to membrane deformation is shown in figure 5 for all cases considered. Consistent with the time-dependent results in figure 4, the cases of $\textit{Ae}=1.7,\hat{\alpha}=35^{o}$} and $\textit{Ae}=1,\hat{\alpha}=55^{o}$} exhibit the largest mean error, $|\Delta\overline{C}_{L_{d}}|\cong 0.44$, with corresponding mean maximum cambers of $19\%$ and $27\%$ of the chord, respectively. As Ae increases, the prediction error rapidly decreases and remains within $|\Delta\overline{C}_{L_{d}}|<0.3$, consistent with the good qualitative agreement observed in figure 4.
In general, low-error cases correspond to scenarios where the time-averaged maximum camber remains below $16\%$ $(\overline{y}_{\max}\leq 0.32)$. An exception is the case of $\textit{Ae}=1.9,\hat{\alpha}=55^{o}$}, which shows good agreement $(\Delta\overline{C}_{L_{d}}=0.21)$ despite a comparatively large camber $(\overline{y}_{\max}=0.45)$. Inspection of the flowfield for this case (figure 4ii.f) indicates that the shear layer reattaches over most of the suction side, resulting in predominantly attached flow. We therefore conclude that the primary requirement for accurate lift prediction using the present theoretical model is attached flow near the trailing edge. The small-camber condition, while related to maintaining attached flow, plays a secondary role.
Figure 5: Absolute value of the time-averaged error in prediction of the lift coefficient due to membrane deformation, $|\Delta\overline{C}_{L_{d}}|$, plotted against (a) aeroelastic number, Ae; (b) mean maximum camber, $\overline{y}_{max}$, for angle-of-attack amplitudes of $\hat{\alpha}=35^{o},55^{o}$. Small Ae values lead to considerable error in $C_{L_{d}}$, where $|\Delta\overline{C}_{L_{d}}|>0.4$, and as Ae increases the error decreases.